OpenStudy (anonymous):

add or subtract the rational expressions, as indicated, and simplify your answer

OpenStudy (anonymous):

\[\frac{ x^2+2x-35 }{ x-5? }+\frac{ x-4 }{ ?x+3 }\]

OpenStudy (tyteen4a03):

Firstly factorize \(x^2 + 2x - 35\), then simplify, and finally continue simplifying.

OpenStudy (anonymous):

\[\frac{ (x+7?) (x-5)}{ (x-5)? }\] so it would look like that right

OpenStudy (tyteen4a03):

No, this is addition not multiplication. :)

OpenStudy (anonymous):

thats why im confused cause thats what my instructor showed us

OpenStudy (anonymous):

Hey what does this ques mark signify???

OpenStudy (anonymous):

@shawilson09

OpenStudy (anonymous):

its automatically put in therei didnt do that